OpenStudy (anonymous):

how do you calculate the angle of a wavelength given wavelength of the light, distance between slits and distance to screen?

OpenStudy (anonymous):

\[\theta = \frac{ \lambda }{ d }\]

OpenStudy (anonymous):

Where d is the distance between the slits.

OpenStudy (hugsnotughs):

I think @IhteshamMalik is correct...do you need help knowing what the symbols mean?

OpenStudy (hugsnotughs):

λ is the wavelength, θ is the angle of deviation of the spectral line, and d is the distance between the slits or the "grating space".

OpenStudy (anonymous):

Constructive interference: \[\sin \theta = \frac{m \lambda}{d}\] where: \(m = 0, \pm1, \pm2, ...\)

OpenStudy (anonymous):

Thanks guys. I'm still confused because how do you know the value for m?

OpenStudy (anonymous):

These are the values I was given. The wavelength of the light is 650. The distance between the slits is 90. The distance to the screen is 400.